# expected utility

### v1

@@ Abstract
This paper studies the equity holders’ mean-variance optimal portfolio choice problem for (non-)protected participating life insurance contracts. We derive explicit formulas for the optimal terminal wealth and the optimal strategy in the multi-dimensional Black-Scholes model, showing the existence of all necessary parameters. In incomplete markets, we state Hamilton-Jacobi-Bellman equations for the value function. Moreover, we provide a numerical analysis of the Black-Scholes market. The equity holders on average increase their investment into the risky asset in bad economic states and decrease their investment over time.
@@ 1 Introduction
This paper aims to provide an explicit formula for the optimal terminal wealth (from the perspective of the equity holders of an insurance company), show that the optimal solution exists (which is non-trivial since the solution includes an additional parameter), and prove an analytical formula for the optimal strategy. Moreover, we derive a Hamilton-Jacobi-Bellman (HJB) equation for incomplete markets and discuss the characteristics of the terminal wealth and the optimal strategy in a numerical analysis.
@@
A significant factor of an optimal policy design is the payoff structure: For a participating life insurance contract, the payoff has at least one point where it is non-differentiable, e.g., at the threshold where the proportional surplus participation starts. (The payoff of the non-protected product also has a second point of non-differentiability at the guarantee value.) At these points, the payoff changes its slope, resulting in a payoff that is generally neither convex nor concave (see Figure 2 for an illustration of the insurer’s payoff). In such situations, one often uses concavification techniques; see, for instance, Larsen [14] or Reichlin [26]. Liang et al. [15] used a generalization of this method. For alternative approaches in non-concave portfolio optimization, see, for instance, Kraft and Steffensen [13], Dai et al. [7], or Qian and Yang [25]. In this work, we use the specific structure of a payoff in a complete market to carefully compare different solution candidates and show that the optimal terminal wealth actually has a closed and, up to the implicit parameters whose existence proof and computation in our setting is delicate, simple form. We also use a standard Lagrangian approach to get the optimal terminal wealth and use the price density process to give the analytic formula for the optimal strategy. This method was used to solve several other optimization problems in complete markets; see, for instance, Basak and Shapiro [2], Cuoco et al. [6], Chen et al. [4], Nguyen and Stadje [23], or Mi et al. [21]. To use this approach for variance, we start to give an equivalent problem in the spirit of Zhou and Li [30]. The ansatz described above then leads to having two multipliers instead of only one Lagrangian multiplier. We obtain a system of non-linear equations with several variables yielding implicit functions whose properties, through various intermediate results, then entail the existence of a solution. Although mean-variance does not respect first-order stochastic dominance, we do get similar results compared to EU-optimization regarding the general form of the optimal solution, but with the terminal wealth having a simpler structure. In particular, in a complete market, the optimal terminal wealth is piecewise linear in the price density. On the other hand, in an incomplete market, the optimal solution itself can only be characterized implicitly as the (viscosity) solution of a certain PDE. Finally, our numerical results show a somewhat different investment behavior compared to EU-maximization. Specifically, we observe that the investment gets more conservative with shorter maturities, and the equity holders on average increase their investment into the risky asset in bad economic states. Moreover, the insurer invests more riskily when offering a non-protected participation life insurance product than when offering a protected one due to a reduced downside potential. Since variance is the most widely used risk measure in industry and the cornerstone of most modern portfolio theory and the finance literature, these results might be potentially relevant from a descriptive point of view.
@@
Section 2 describes participating life insurance contracts and briefly introduces the functional to optimize. In Section 3, we explicitly show the optimal terminal wealth, the optimal strategy, and the existence of the necessary parameters in the Black-Scholes model. In Section 4, we give the SDEs for possibly incomplete markets, and Section 5 analyzes some numerical results for the Black-Scholes model. Finally, Section 6 concludes the paper.
@@ 2.2 Optimization Functional
Before defining the optimization functional, let us introduce the basic financial market, which in Section 3 and 5 will become the Black-Scholes model.
@@ 3.1 Derivation of the optimal terminal wealth
    Note that the representation of $\hat{X}_{T}$ also holds in a general complete market (without necessarily using a Black-Scholes market model) if we assume the existence of the variables $y$ and $\lambda$. However, that these variables exist is then not clear.
@@
    The following paragraph considers $L$ as a function of $X$. Obviously, $L$ is not smooth in $k_{1}$ and $k_{2}$, but continuous in $[0,\infty)$. Hence, we optimize $L$ in the regions $[0,k_{1}]$, $[k_{1},k_{2}]$, and $[k_{2},\infty)$ separately and compare afterwards the minimal points. Now, for the individual optimization, we get:
@@ 3.2 Derivation of the optimal strategy
4 Optimization in a possibly incomplete market
In this subsection, we consider a financial market without the completeness assumption such that we can write the wealth process as:
$\mathrm{d}X_{t}=b(X_{t},u_{t})\mathrm{d}t+\sigma(X_{t},u_{t})\mathrm{d}W_{t},$
where $W$ is again the $d$-dimensional Brownian Motion generating the filtration $\mathcal{F}$, and $b$ and $\sigma$ are measurable functions which satisfy a uniform Lipschitz condition.
First, Lemma 3.1 also holds in incomplete markets since we do not use completeness in its proof. Now, for this optimization, we denote for $(t,x)\in[0,T]\times\mathbb{R}$ the optimal value functional by $V$:
$V(t,x)=\sup_{u\in\mathcal{U}(t,x)}\tilde{J}(t,T,u,x)=\sup_{u\in\mathcal{U}(t,x)}\mathbb{E}[\lambda F(0,T,u,x_{0})-\gamma F(0,T,u,x_{0})^{2}],$
where $\mathcal{U}(t,x)$ denotes the subset of $\mathcal{U}$ with processes starting at $t$ and $X_{t}=x$. Remember that $\lambda=1+2\gamma\mathbb{E}\left[F(0,T,\hat{u},x_{0})\right]$, i.e., $\lambda$ depends on the optimal strategy $\hat{u}$.
Theorem 4.1.
If $V\in C^{1,2}$, then for every $\lambda\geq 0$ the optimal value functional $V$ is the solution of the following SDE for all $(t,x)\in[0,T)\times\mathbb{R}$:
$-\displaystyle\frac{\mathrm{d}V}{\mathrm{d}t}(t,x)-\sup_{u\in\mathcal{U}}\mathcal{L}^{u}V(t,x)$ $=0,$
$V(T,x)$ $=F(T,T,0,x),$
where the operator $\mathcal{L}^{u}$ is defined as
$\mathcal{L}^{u}v(t,x)$ $:=b(x,u)\displaystyle\frac{\mathrm{d}V}{\mathrm{d}x}(t,x)+\frac{1}{2}\sigma(x,u)\sigma^{T}(x,u)\displaystyle\frac{\mathrm{d}^{2}V}{\mathrm{d}x^{2}}(t,x),$
where tr denotes the trace of a matrix.
Proof. Obviously, $\tilde{J}$ satisfies a quadratic growth condition. Then, the result follows from Chapter 3 in Pham [24].
The regularity assumption of $V$ is typically challenging to prove. Therefore, we also give the result for viscosity solutions where this assumption is not needed:
Theorem 4.2.
Let the control space $\mathcal{U}$ be compact and the Hamiltonian $H$ be defined as usual, i.e.:
$H$ $:[0,T)\times\mathbb{R}\times\mathbb{R}\times\mathbb{R}\to\mathbb{R},$
$H(t,x,p,M)$ $:=\sup_{u\in\mathcal{U}}\left[b(x,u)\displaystyle\frac{\mathrm{d}V}{\mathrm{d}x}(t,x)+\frac{1}{2}\sigma(x,u)\sigma^{T}(x,u)\displaystyle\frac{\mathrm{d}^{2}V}{\mathrm{d}x^{2}}(t,x)\right].$
If $V$ is locally bounded on $[0,T)\times\mathbb{R}$, then for every $\lambda\geq 0$, $V$ is a viscosity solution of the following HJB equation for $(t,x)\in[0,T)\times\mathbb{R}$:
$-\displaystyle\frac{\mathrm{d}V}{\mathrm{d}t}(t,x)-H\left(t,x,\displaystyle\frac{\mathrm{d}V}{\mathrm{d}x}(t,x),\displaystyle\frac{\mathrm{d}^{2}V}{\mathrm{d}x^{2}}(t,x)\right)$ $=0,$
$V(T,x)$ $=F(T,T,0,x).$
Proof. The result follows from Chapter 4 in Pham [24].
Note that it is possible to relax the assumption of $\mathcal{U}$ being compact, leading to a possible singularity in the Hamiltonian. This leads to a complicated structure of the HJB equation and the terminal value condition.
@@ 5 Numerical Results
For the participating life insurance contract, we use the parameters $k_{0}=0$, $k_{1}=2.5$, $k_{2}=7$, $\alpha=1$, and $\alpha_{2}=0.25$ for the non-protected insurance product. We change the values of $k_{0}$ and $k_{1}$, i.e., $k_{0}=2.5$ and $k_{1}=0$, to get the protected product. The initial wealth, i.e., the sum of the premiums and the initial capital from the insurance company, is given by $x_{0}=4$, the time horizon by $T=10$, and the risk aversion by $\gamma=0.25$. We assume a constant risk-free interest rate of $r:=r_{t}\equiv 0.02$ for the financial market and consider one risky asset with constant mean $\mu:=\mu_{t}\equiv 0.08$ and volatility $\sigma_{t}\equiv 0.2$. In particular, the Sharpe ratio is also constant and given by $\kappa_{t}\equiv 0.3$. For the implementation, we used a time step of $\delta=0.01$, i.e., we have 1000 intermediate points. Moreover, we simulate 1000 realizations of the Brownian Motion.
@@
In Figure 6, we show the optimal wealth process and the optimal share into the risky asset over time for the non-protected participating life insurance product. In both figures, the red line shows the average of the 1000 realizations, and each of the ten black lines shows a single realization. Note that for the optimal strategy, we used a weighted average with the weight given by the absolute investment amount (which also holds for all other averages in this section). Furthermore, we get $\lambda\approx 3.423$ and $y\approx 0.860$. The average optimal wealth process develops from $4$ to approximately $7.8$, where the increase is slightly larger in the beginning than in the end. A terminal wealth of about $7.8$ corresponds to a terminal wealth for the insurer of $5.1$ and a wealth of $2.7$ for the policyholder. Note that this is higher than the guarantee value, which is $2.5$. When exercising the optimal strategy, we start with high investment into the risky asset (approximately $110\%$ of our wealth), which decreases relatively constant over time. The final optimal investment share into the risky asset is around $40\%$. When analyzing the strategies in more detail, one observes that the most risky investments are taken when the economy has poorly evolved until then, i.e., when $\xi_{t}$ is high, while the least risky investments are taken in cases where the economy has developed well. These results are reasonable since when the wealth is below the guarantee, the insurer has nothing to lose anymore, while the wealth is already over the second threshold $k_{2}$ (i.e., the policyholder also participates in the surplus over $k_{2}$), the upside potential of the insurer is reduced, whereas the downside potential is not (or only slightly).
@@
In Figure 7, we show the optimal wealth process and the optimal share into the risky asset over time for the protected participating life insurance product. Again, the red line shows the average of the 1000 realizations, and each of the ten black lines shows a single realization. Moreover, the variables $\lambda$ and $y$ are given by $\lambda\approx 2.893$ and $y\approx 1.003$. In this case, the optimal value evolves only to around $6.8$, corresponding to a terminal wealth of the insurance company of $4.3$ since we are below $k_{2}$, i.e., there is no surplus participation of the policyholder. The optimal strategy starts risky but not as risky as in the non-protected case, with around $75\%$ of the wealth invested into the risky asset. It also decreases over time, but the reduction is smaller, and the final optimal investment percentage is at around $30\%$ of the wealth. As in the non-protected case, the realizations in the best economic states correspond with the least risky investments and vice versa. One of the shown strategies is rather inconspicuous over most of the time but has a major change close to the final time point $T=10$, i.e., a drop-down from $u_{9.93}=0.2$ to $u_{9.99}=0.00001$. Such an effect (also in the other way, i.e., an upward move) happens for several realizations when their wealth for time points $t$, which are close to the maturity $T$, is close to the second threshold $k_{2}$ since the optimal terminal wealth as a function of $\xi$ has a small plateau at $k_{2}$, i.e., if $\xi_{t}\in(\tilde{\alpha}\hat{\xi},\xi_{2}^{*}]$. Hence, for $\xi_{t}\approx\tilde{\alpha}\hat{\xi}$ (resp. $\xi_{t}\approx\tilde{\alpha}\xi_{2}^{*}$) for $t$ close to $T$, the upside (resp. downside) potential is reduced and the optimal strategy is to invest safely (resp. risky).
@@
In Figure 8, we compare the optimal wealth before splitting it between the policyholders and the insurance company and the optimal strategy of our two participating life insurance products with mean-variance to a non-participating investment, i.e., we set $\alpha_{2}=0$, $k_{0}=0$, and $k_{1}=0$. Moreover, we set the initial wealth such that the mean of $\hat{X}_{T}$ is approximately equal to $7.787$. Therefore, we get the following initial values: $x_{0}={\scriptsize\begin{cases}4&\text{ non-protected product},\\ 4.765&\text{ protected product},\\ 4.738&\text{ no participation}.\end{cases}}$ The lines show again an average of $1000$ realizations each. The figure shows that the investor makes the riskiest investment when offering a non-protected insurance product and approximately identical investments when offering one of the other two products. When offering the protected product compared to the non-participation product, it is remarkable that the investor invests slightly safer in the beginning and somewhat riskier close to maturity. This investment behavior leads to the highest wealth gain for the non-protected product since, on average, the risky asset performs better than the risk-less asset (due to $\mu>r$). For the other two products, the wealth gain is approximately equal. The protected life insurance investment strategy is relatively close to the non-participating strategy since a fixed-guarantee payment only minimally influences the optimal strategy. Then, the only remaining difference in the payoff shape is when the wealth is above $k_{2}$, which happens in our chosen parametrization either relatively late or not at all (see Figure 7). The payoff structure for the non-protected insurance product differs highly in the values due to the reduced downside potential below $k_{1}$.
@@
In Figure 9, we compare the optimal wealth and strategy for the non-protected and the protected life insurance contract when considering mean-variance resp. EU-optimization. We take the result for optimizing EU from Lin et al. [16] as in Figure 4 with parameter values $\tilde{\gamma}=0.125$ and $\tilde{\lambda}=2$. As in the other figures, the lines show the average of $1000$ realizations, and we state the total wealth (before splitting it between the policyholders and the insurance company). As in the previous figure, we again choose the initial value such that the terminal wealth is approximately equal for all products, i.e., we take $x_{0}={\scriptsize\begin{cases}4&\text{ non-protected product},\\ 4.765&\text{ protected product}\end{cases}}$ for the mean-variance optimization and $x_{0}={\scriptsize\begin{cases}3.343&\text{ non-protected product},\\ 3.616&\text{ protected product}\end{cases}}$ for the EU-optimization. Both optimizations have in common that the insurance company offering a non-protected product invests riskier than the insurance company offering the protected product due to the reduced downside potential. When comparing the strategies, we observe that the strategies optimizing EU become riskier over time (except for the final time points). In contrast, the strategies optimizing mean-variance get less risky over time, as already discussed. Since we chose the two (not comparable) risk aversion factors such that the initial investments in the risky asset are similar, the wealth gain of the EU-optimization is higher. The decreasing portion invested into the risky asset is typical for a pre-commitment mean-variance optimization strategy (this can be seen by implementing the optimal mean-variance strategy by Zhou and Li [30] when using typical ranges for the parameters), which is structurally different from EU-optimization strategies.
@@
Figure 9: Comparison of the optimal wealth (left) and the optimal strategy (right) between a (non-)protected participating life insurance contract optimizing mean-variance and EU.

### v2

@@ Abstract
This paper studies the equity holders’ mean-variance optimal portfolio choice problem for (non-)protected participating life insurance contracts. We derive explicit formulas for the optimal terminal wealth and the optimal strategy in the multi-dimensional Black-Scholes model, showing the existence of all necessary parameters. Moreover, we provide a numerical analysis of the Black-Scholes market. The equity holders on average increase their investment into the risky asset in bad economic states and decrease their investment over time.
@@ 1 Introduction
This paper aims to provide an explicit formula for the optimal terminal wealth (from the perspective of the equity holders of an insurance company), show that the optimal solution exists (which is non-trivial since the solution includes an additional parameter), and prove an analytical formula for the optimal strategy. Moreover, we discuss the characteristics of the terminal wealth and the optimal strategy in a numerical analysis.
@@
A significant factor of an optimal policy design is the payoff structure: For a participating life insurance contract, the payoff has at least one point where it is non-differentiable, e.g., at the threshold where the proportional surplus participation starts. (The payoff of the non-protected product also has a second point of non-differentiability at the guarantee value.) At these points, the payoff changes its slope, resulting in a payoff that is generally neither convex nor concave (see Figure 2 for an illustration of the insurer’s payoff). In such situations, one often uses concavification techniques; see, for instance, Larsen [22] or Reichlin [35]. Liang et al. [24] used a generalization of this method. For alternative approaches in non-concave portfolio optimization, see, for instance, Kraft and Steffensen [21], Dai et al. [14], or Qian and Yang [34]. In this work, we use the specific structure of a payoff in a complete market to carefully compare different solution candidates and show that the optimal terminal wealth actually has a closed and, up to the implicit parameters whose existence proof and computation in our setting is delicate, simple form. We also use a standard Lagrangian approach to get the optimal terminal wealth and use the price density process to give the analytic formula for the optimal strategy. This method was used to solve several other optimization problems in complete markets; see, for instance, Basak and Shapiro [4], Cuoco et al. [13], Chen et al. [8], Chen et al. [7], Nguyen and Stadje [33], Mi et al. [31], Chen et al. [10], or Avanzi et al. [1]. To use this Lagrangian approach for mean-variance, we start to give an equivalent problem in the spirit of Zhou and Li [39]. The ansatz described above then leads to having two multipliers instead of only one Lagrangian multiplier. We obtain a system of non-linear equations with several variables yielding implicit functions whose properties, through various intermediate results, then entail the existence of a solution. Although mean-variance does not respect first-order stochastic dominance, we do get similar results compared to EU-optimization regarding the general form of the optimal solution, but with the terminal wealth having a simpler structure. In particular, in a complete market, the optimal terminal wealth is piecewise linear in the price density. Compared to the papers discussed above, the main difference of our ansatz is to consider an optimization functional, where we include an option-like payoff of the participation instead of adding further conditions to a non-option-like payoff. We show in Proposition 3.6 that our optimization corresponds to a target problem, where one aims to be as close as possible to certain exogenously given constants. This alternative formulation is close to Avanzi et al. [1], who minimize a variance-like function under solvency conditions. However, while they face a 1-target problem our optimization corresponds to a 2-target problem, where the exogenously given target depends on the region of the terminal wealth. An additional constraint ensures that mass can only be shifted from one region to another at a rate which depends on the amount of participation. Target-based optimization is common in the literature as an alternative to mean-variance, see for instance Avanzi et al. [1], Menoncin and Vigna [28], or Li and Forsyth [23]. Finally, our numerical results show a somewhat different investment behavior compared to EU-maximization, in particular for S-shaped utility functions. Specifically, we observe that the investment gets more conservative with shorter maturities, and the equity holders on average increase their investment into the risky asset in bad economic states. Moreover, the insurer invests more riskily when offering a non-protected participation life insurance product than when offering a protected one due to a reduced downside potential. Since variance is the most widely used risk measure in industry and the cornerstone of most modern portfolio theory and the finance literature, these results might be potentially relevant from a descriptive point of view.
@@
Section 2 describes participating life insurance contracts and briefly introduces the functional to optimize. In Section 3, we explicitly show the optimal terminal wealth, the optimal strategy, and the existence of the necessary parameters in the Black-Scholes model. Section 4 analyzes some numerical results for the Black-Scholes model. Finally, Section 5 concludes the paper.
@@ 2.2 Optimization Functional
Before defining the optimization functional, let us introduce the basic financial market, which in Section 3 and 4 will become the Black-Scholes model.
@@
Remark 2.1.
Since the problem in this paper is a variant of a mean-variance optimization, it inherits also the time-inconsistency of the classical mean-variance approach, see, for instance, Basak and Chabakauri [3] or Björk and Murgoci [5]. There are several ways to approach this time-inconsistency. We take in this paper the perspective of a so-called pre-committed investor. Such an investor optimizes once in the beginning and sticks to this strategy, i.e., the chosen strategy of the investor is not dynamically reoptimized. Note that due to the function $F$, our problem is not covered by the general theory from Björk and Murgoci [5] which assumes a sophisticated investor, i.e., she/he knows that her/his optimal strategy changes over time and only optimizes over all strategies she/he would actually follow.
@@ 3.1 Derivation of the optimal terminal wealth
    The representation of $\hat{X}_{T}$ also holds in a general complete market (without necessarily using a Black-Scholes market model) if we assume the existence of the multipliers $y$ and $\lambda$. However, that these exist is then not clear. Note that the price density process, $\xi$, in a different complete market would change, but still admit a geometric form. For example, if we replace the Brownian Motion $W$ by a Poisson process $N$ with parameter $\lambda_{2}$ (in the Black-Scholes model), the price density admits a dynamic of the form (see, e.g., Cont and Tankov [12, p.303]):
    $\mathrm{d}\xi_{t-}$ $=\xi_{t-}(\lambda_{2}-\lambda_{1}-r_{t})\mathrm{d}t-\xi_{t-}\ln\tfrac{\lambda_{2}}{\lambda_{1}}\mathrm{d}N_{t},$
    where $\lambda_{1}>0$ is an appropriately chosen constant.
@@
In the following proposition, we give an alternative formulation for our optimization problem. Again, we give the proof in Appendix A.
Proposition 3.6.
Let $\tilde{L}(x)={\scriptsize\begin{cases}\gamma\alpha^{2}k_{0}^{2},&\text{if }x<k_{1},\\ \gamma\alpha^{2}(x-k_{1}-k_{0})^{2},&\text{if }k_{1}\leq x<k_{2},\\ \gamma\tilde{\alpha}^{2}(x-k_{2}+\tfrac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0}))^{2},&\text{if }x\geq k_{2}.\end{cases}}$ and consider the optimization problem $\min_{u\in\mathcal{U}}\mathbb{E}[\tilde{L}(X_{T})]$ under the constraint $\mathbb{E}[\alpha(X_{T}(\lambda)-k_{1})\mathbbm{1}_{k_{1}\leq X_{T}(\lambda)<k_{2}}+\linebreak\tilde{\alpha}(X_{T}(\lambda)-k_{2}+\tfrac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}))\mathbbm{1}_{X_{T}(\lambda)\geq k_{2}}]=\alpha k_{0}$ and the budget constraint $\mathbb{E}[\xi_{T}X_{T}(y)]=\xi_{0}x_{0}$, where $\lambda$ and $y$ are the respective Lagrangian multipliers.
Then, the optimal terminal wealth is identical to the one of our main problem, i.e., given by (3.4).
Note that the loss function $\tilde{L}$ is continuous and non-decreasing in $x$, and that the definition and therefore the interpretation of the parameter $\lambda$ is different than for mean-variance. However, Proposition 3.6 still gives an equivalent formulation. As already stated in the introduction, this formulation is a 2-target problem. Depending on the region of the terminal wealth, we want to minimize the distance to different exogenously given target values ($k_{1}+k_{0}$ if the terminal wealth $X\in[k_{1},k_{2})$ respectively to $k_{2}-\tfrac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0})$ if $X\geq k_{2}$). Moreover, the first constraint entails that one can only reduce mass in one region when one simultaneously increases the mass in another region where the reduction and the enlargement are, in general, not of equal size. Especially, the weight of the region $[k_{2},\infty)$ is heavily influenced by the participation rate $\alpha_{2}$ through the pre-factor $\tilde{\alpha}(=\alpha-\alpha_{2})$.
@@
    The following paragraph considers $L$ as a function of $X$. Obviously, $L$ is not smooth in $k_{1}$ and $k_{2}$, but it is continuous in $[0,\infty)$. Hence, we optimize $L$ in the regions $[0,k_{1}]$, $[k_{1},k_{2}]$, and $[k_{2},\infty)$ separately and compare afterwards the minimal points. Now, for the individual optimization, we get:
@@ 4 Numerical Results
For the participating life insurance contract, we use the parameters $k_{0}=0$, $k_{1}=2.5$, $k_{2}=7$, $\alpha=1$, and $\alpha_{2}=0.25$ for the non-protected insurance product. We change the values of $k_{0}$ and $k_{1}$, i.e., $k_{0}=2.5$ and $k_{1}=0$, to get the protected product. The initial wealth, i.e., the sum of the premiums and the initial capital from the insurance company, is given by $x_{0}=4$, the time horizon by $T=10$, and the risk aversion by $\gamma=0.25$. We assume a constant risk-free interest rate of $r:=r_{t}\equiv 0.02$ for the financial market and consider one risky asset with constant mean $\mu:=\mu_{t}\equiv 0.08$ and volatility $\sigma_{t}\equiv 0.2$. In particular, the Sharpe ratio is also constant and given by $\kappa_{t}\equiv 0.3$. For the implementation, we used a time step of $\delta=0.01$, i.e., we have 1000 intermediate points. Moreover, we simulate 10000 realizations of the Brownian Motion.
@@
In both, Figure 4 and Figure 5, we can observe a jump of the optimal terminal wealth in the non-protected case, which is typical for problems with option-like constraints, see for instance Basak and Shapiro [4], Chen et al. [8], Nguyen and Stadje [33], Chen et al. [9], or Kraft and Steffensen [21]. The reason is that for certain outcomes, the insurance company’s incentive change and therefore it can gain additional utility by changing to a different regime. [^6] For instance, if the entire wealth for certain outcomes is transferred to the policyholders, the insurance company will target a terminal wealth of zero in these states in order to gain additional wealth in states where it benefits more. (Note that the overall expectation is fixed due to the budget constraint.) Moreover, the piecewise linearity observed also occurs in, e.g., Avanzi et al. [1] who considered a minimization of a quadratic function under solvency constraints including option-like constraints as the expected shortfall. This piecewise linearity is typical for problems with some kind of quadratic penalties, whereas the option-like constraints entail the existence of the plateaus and the jumps.
In Figure 6, we show the optimal wealth process and the optimal share into the risky asset over time for the non-protected participating life insurance product. In both figures, the red line shows the average of the $10000$ realizations, and each of the ten black lines shows a single realization. Note that for the optimal strategy, we used a weighted average with the weight given by the absolute investment amount (which also holds for all other averages in this section). Furthermore, we get $\lambda\approx 3.423$ and $y\approx 0.860$. The average optimal wealth process develops from $4$ to approximately $7.6$, where the increase is slightly larger in the beginning than in the end. A terminal wealth of about $7.6$ corresponds to a terminal wealth for the insurer of $4.95$ and a wealth of $2.65$ for the policyholder. Note that this is higher than the guarantee value, which is $2.5$. When exercising the optimal strategy, we start with high investment into the risky asset (approximately $110\%$ of our wealth), which decreases relatively constant over time. The final optimal investment share into the risky asset is around ${45}\%$. When analyzing the strategies in more detail, one observes that the most risky investments are taken when the economy has poorly evolved until then, i.e., when $\xi_{t}$ is high, while the least risky investments are taken in cases where the economy has developed well. These results are reasonable since when the wealth is below the guarantee, the insurer has nothing to lose anymore, while the wealth is already over the second threshold $k_{2}$ (i.e., the policyholder also participates in the surplus over $k_{2}$), the upside potential of the insurer is reduced, whereas the downside potential is not (or only slightly).
@@
In Figure 7, we show the optimal wealth process and the optimal share into the risky asset over time for the protected participating life insurance product. Again, the red line shows the average of the $10000$ realizations, and each of the ten black lines shows a single realization. Moreover, the variables $\lambda$ and $y$ are given by $\lambda\approx 2.893$ and $y\approx 1.003$. In this case, the optimal value evolves only to around $6.7$, corresponding to a terminal wealth of the insurance company of $4.2$ since we are below $k_{2}$, i.e., there is no surplus participation of the policyholder. The optimal strategy starts risky but not as risky as in the non-protected case, with around $75\%$ of the wealth invested into the risky asset. It also decreases over time, but the reduction is smaller, and the final optimal investment percentage is at around $35\%$ of the wealth. As in the non-protected case, the realizations in the best economic states correspond with the least risky investments and vice versa. One of the shown strategies is rather inconspicuous over most of the time but has a major change close to the final time point $T=10$, i.e., a drop-down from $u_{9.93}=0.2$ to $u_{9.99}=0.00001$. Such an effect (also in the other way, i.e., an upward move) happens for several realizations when their wealth for time points $t$, which are close to the maturity $T$, is close to the second threshold $k_{2}$ since the optimal terminal wealth as a function of $\xi$ has a small plateau at $k_{2}$, i.e., if $\xi_{t}\in(\tilde{\alpha}\hat{\xi},\xi_{2}^{*}]$. Hence, for $\xi_{t}\approx\tilde{\alpha}\hat{\xi}$ (resp. $\xi_{t}\approx\tilde{\alpha}\xi_{2}^{*}$) for $t$ close to $T$, the upside (resp. downside) potential is reduced and the optimal strategy is to invest safely (resp. risky).
@@
In Figure 8, we compare the optimal wealth before splitting it between the policyholders and the insurance company and the optimal strategy of our two participating life insurance products with mean-variance to a non-participating investment, i.e., we set $\alpha_{2}=0$, $k_{0}=0$, and $k_{1}=0$. Moreover, we set the initial wealth such that the mean of $\hat{X}_{T}$ is approximately equal to $7.59$. Therefore, we get the following initial values: $x_{0}={\scriptsize\begin{cases}4&\text{ non-protected product},\\ {4.692}&\text{ protected product},\\ {4.665}&\text{ no participation}.\end{cases}}$ The lines show again an average of $10000$ realizations each. The figure shows that the investor makes the riskiest investment when offering a non-protected insurance product and approximately identical investments when offering one of the other two products. When offering the protected product compared to the non-participation product, it is remarkable that the investor invests slightly safer in the beginning and somewhat riskier close to maturity. This investment behavior leads to the highest wealth gain for the non-protected product since, on average, the risky asset performs better than the risk-less asset (due to $\mu>r$). For the other two products, the wealth gain is approximately equal. The protected life insurance investment strategy is relatively close to the non-participating strategy since a fixed-guarantee payment only minimally influences the optimal strategy. Then, the only remaining difference in the payoff shape is when the wealth is above $k_{2}$, which happens in our chosen parametrization either relatively late or not at all (see Figure 7). The payoff structure for the non-protected insurance product differs highly in the values due to the reduced downside potential below $k_{1}$.
@@
In Figure 9, we compare the optimal wealth and strategy for the non-protected and the protected life insurance contract when considering mean-variance resp. EU-optimization. We take the result for optimizing EU from Lin et al. [25] as in Figure 4 with parameter values $\tilde{\gamma}=0.125$ and $\tilde{\lambda}=2$. As in the other figures, the lines show the average of $10000$ realizations, and we state the total wealth (before splitting it between the policyholders and the insurance company). As in the previous figure, we again choose the initial value such that the terminal wealth is approximately equal for all products, i.e., we take $x_{0}={\scriptsize\begin{cases}4&\text{ non-protected product},\\ {4.692}&\text{ protected product}\end{cases}}$ for the mean-variance optimization and $x_{0}={\scriptsize\begin{cases}{3.31}&\text{ non-protected product},\\ {3.59}&\text{ protected product}\end{cases}}$ for the EU-optimization. Both optimizations have in common that the insurance company offering a non-protected product invests riskier than the insurance company offering the protected product due to the reduced downside potential. When comparing the strategies, we observe that the strategies optimizing EU become riskier over time (except for the final time points). In contrast, the strategies optimizing mean-variance get less risky over time, as already discussed. Since we chose the two (not comparable) risk aversion factors such that the initial investments in the risky asset are similar, the wealth gain of the EU-optimization is higher. The decreasing portion invested into the risky asset is typical for a pre-commitment mean-variance optimization strategy (this can be seen by implementing the optimal mean-variance strategy by Zhou and Li [39] when using typical ranges for the parameters), which is structurally different from EU-optimization strategies.
@@
Figure 9: Comparison of the optimal wealth (left) and the optimal strategy (right) between a (non-)protected participating life insurance contract optimizing mean-variance and EU with S-shaped utility from Lin et al. [25].
In Figure 10, we show another comparison for the optimal wealth and strategy for the non-protected and the protected life insurance contract when considering mean-variance resp. EU-optimization. Here, we take for the EU-optimization an exponential utility function $U(x)=1-e^{-\hat{\gamma}x}$, where $\hat{\gamma}=0.5$ denotes the risk aversion of the investor. Note that of course this risk aversion parameter is not directly comparable to the risk aversion parameter from our mean-variance model. (For small risks though $U$ is actually related to mean-variance by a Taylor-expansion.) The optimal wealth and strategy are calculated based on Korn and Trautmann [20]. Due to their assumptions, we change the risk-free interest rate to $r=0$. Again, we show the average of $10000$ realizations, state the total wealth (before splitting it between the policyholders and the insurance company), and choose the initial value such that the terminal wealth is approximately equal for all products, i.e., we take $x_{0}={\scriptsize\begin{cases}4&\text{ non-protected product},\\ 4.77&\text{ protected product}\end{cases}}$ for the mean-variance optimization and $x_{0}={\scriptsize\begin{cases}6.224&\text{ non-protected product},\\ 6.262&\text{ protected product}\end{cases}}$ for the EU-optimization. We obtain again that the insurance company offering a non-protected product invests riskier than the insurance company offering the protected product due to the reduced downside potential, where the difference between the strategies is larger for the mean-variance optimization. In particular, we observe for exponential EU-optimization that the optimal strategy is relatively similar for the non-protected and the protected life insurance products. Furthermore, these strategies are decreasing for longer maturities and slightly increasing for short maturities, and are closer to the mean-variance optimization than the ones discussed in Figure 9 for S-shaped utility. However, the change in the strategies for exponential EU-maximization over time is less than for mean-variance optimization.
Figure 10: Comparison of the optimal wealth (left) and the optimal strategy (right) between a (non-)protected participating life insurance contract optimizing mean-variance, or EU with an exponential utility function.